[Hallc_running] [New Logentry] RC daily update

billlee at jlab.org billlee at jlab.org
Fri Sep 14 09:25:01 EDT 2018

Logentry Text:
>From the MCC meeting this morning:
--> Both linacs will be pumped down on Monday (Sep/17). North Linac partial cool down has started and is going well.

Hall C update:
--> All systems remain shut.
--> The HMS and SHMS motors were installed. The SHMS motor was electrically connected yesterday and the HMS motor is expected to be connected this morning.
--> Hall lock down is expected by the end of the day-shift next Thursday (4pm, Sep/20th). 
--> Shifts remain canceled until 8am, Sep/20th (Next Thursday).
--> We will not have a RC meeting today and weekend. 

Planned work for next week:
--> Howard will restore DC normal gas flow on Monday (Sep/17), DC will be ready for operation by Tuesday noon. 
--> Steve and Joe will test the newly install motors next Monday (Sep/17).
--> Beam permit check list will be performed early next week.
--> System will be turned on next Monday, and a cosmic ray test is planned for next Tuesday (Sep/18) afternoon.


This is a plain text email for clients that cannot display HTML.  The full logentry can be found online at https://logbooks.jlab.org/entry/3594475
-------------- next part --------------
An HTML attachment was scrubbed...
URL: <https://mailman.jlab.org/pipermail/hallc_running/attachments/20180914/29e56dea/attachment.html>

More information about the Hallc_running mailing list